OpenStudy (anonymous):

the sum of 4 times a number and 14 is 16. What is the number

OpenStudy (hoblos):

4x+14 =16 4x=2 x=2/4=1/2

OpenStudy (anonymous):

Let a number = x 4x+14=16 4x=2 x=1/2